OpenStudy (anonymous):

is the series sum[n=1,infinity](1-sqrt(1-(1/n))) convergent or divergent?

OpenStudy (anonymous):

Would you be willing to try the integral test?

OpenStudy (anonymous):

@melaniebrown, you're given a positive and decreasing series, do you think you'd be able to integrate \[\int_1^\infty \left(1-\sqrt{1-\frac{1}{x}}\right)~dx ~~?\]

OpenStudy (anonymous):

Alternatively, you can try comparing to a known divergent/convergent series and see what you get. I have a feeling this series diverges, so I would try comparing to \(\sum\dfrac{1}{\sqrt n}\).